OpenStudy (anonymous):

if tan(A+B) : tan(A-B) = 1 : 2 and Sin 2A + K Sin 2B = 0 , then find the Value of K ?

OpenStudy (anonymous):

@waterineyes @satellite73

OpenStudy (anonymous):

*

OpenStudy (anonymous):

@shubhamsrg @hartnn

OpenStudy (shubhamsrg):

sin(a+b) cos(a-b) / sin(a-b) cos(a+b) = 1/2 =>sin(2a) + sin(2b) / sin(2a) - sin(2b) = 1/2 we know, sin2a /sin2b = -k on little observation/substitution , we see we have -k +1 / -k -1 = 1/2 just solve for k

OpenStudy (anonymous):

thxxxx

OpenStudy (shubhamsrg):

glad to help! :)